OpenStudy (anonymous):

x/(x+5)(x+3) minus 9/(x+3)(x-3)

OpenStudy (anonymous):

1) (x - 5)/3x < 3 x - 5 < 9x 8x > - 5 x > - 5/8 3) (x + 5) /4x > 3 x + 5 > 12x 11x < 5 x < 5/11

OpenStudy (anonymous):

If you need to combine them multiply the first term by (x-3) / (x-3) and the second term by (x+5) / (x+5). then you have common denominators.

jhonyy9 (jhonyy9):

x(x-3)-9(x+5)=x2-3x-9x-45=x2 -12x -45 so than will be x2 -12x -45 -------------- (x+5)(x-3)(x+3)